The displacement field in a steel machine component (Young’s modulus, E = 210 GPa and Poisson’s
ratio, ? = 0.33) is given (in metres) by the following expressions:
ux = c(x2 + ?) + 2c
uy = 2cx2 − cy
uz = c? + 5c?
where c = 10−4.
For a point (x, y, z) positioned at (1, 1, 1) calculate
a) The strain components; [4 marks]
b) The stress components and sketch a stress element showing how they act; [4 marks]
c) The body force needed to maintain a state of equilibrium. [4 marks]
Answer #1


Solution Given data- E=2100Pa = 210x103MPa 2- 0.33 c=104 Ux=((x+y)+20 Uy= 2 cx²-cy Uz = cy+5cz. point. (161). (6) The stres s2 Ку $ук оу dux dy dx 4 = c(+y) +2 у = 24х2 су Uzy+cz = C + 2. С. 27 (1 + 4 ) (144x) е - Sc 25x104 c Хуу = 1+x) 2. x 169 8xy2 . Strain components 2 2х10 9 2.5x10 4 2.5 xo-9 -10-4 0.5000s o 5x10-4 0.5X104 2. ?: O 2. - | 0.5x304 2 of 5 (6) Noc, calcul2= CH (1-2) (It) (1-27) 0-33X210X103 (10.33 (: - 2x0.33) 0.33X2108103 1.33 X 0.34 - 1532 50.77 MPa Txx = 4+26) Exxt ) Eyy t X022= 16xx + 16уу + (1+24) erz = 15 - 3454/* x 3х1 5 15 345 х/05x (-1) -- 15 34 SP-+ 3x 7 8 94x703) x 510-9 30:45 15. 3 3 1 1557 १८ xhe का Kayz >OK 12 247 2 ANU 2106 Now, for state of equilibrium- 2 Pxx + تها 2xy 1 / 2 2x2 + Bx = 0 ův We know €2250 Exx=2(x, x x = (1+29) Exx + x tyy+X EzzSimulary in 7-direction & tay + y Bay + 22 2yz + By = 0. 1 / 2xy = x 3:47 1+48) < af? from 0] 46.4 = 4x10-4x4 2 Y2 = 0. 22 2

i try my best to write in a neat and clean manner.

if u got it plz upvote.